OpenStudy (anonymous):

Factor this pellet :| 3(a-4)^5 (2a +3)^-2 -1/2(a-4)^4 (2a+3)^-1

OpenStudy (pfenn1):

Is this your expression?\[\frac{3(a-4)^5 (2a +3)^{-2} -1}{2(a-4)^4 (2a+3)^{-1}}\]

OpenStudy (anonymous):

\[3(a-4)^{5}(2a+3)^{-2}-1/2(a-4)^{4}(2a+3)^{-1}\]

OpenStudy (anonymous):

Thats my expression

OpenStudy (pfenn1):

Remember\[x^{-a}=\frac{1}{x^a}\]and \[x^ax^b=x^{a+b}\]

OpenStudy (anonymous):

this is the textbook step that I don't get:

OpenStudy (precal):

factor out (a-4)^4 and (2a+3)^-1

OpenStudy (anonymous):

\[1/2(a-4)^{4}(2a+3)^{-2} [6(a-4) -(2a+3)]\]

OpenStudy (anonymous):

ohh ok

OpenStudy (precal):

did you understand my instructions? pfenn1 thought it was another type of problem

OpenStudy (anonymous):

yeah but shouldn't it be (2a+3)^-1 that gets factored out

OpenStudy (anonymous):

ok lol you said that but in the book, it has

OpenStudy (anonymous):

(2a+3)^-2 is factored out :S

OpenStudy (precal):

follow the book, I forget if it is -1 or -2

OpenStudy (anonymous):

but it should be -1 because then: -1 - -2 = 1 which shut leave you with (2a+3)

OpenStudy (anonymous):

but they got (2a+3) by factoring out (2a+3)^-2

OpenStudy (anonymous):

from (2a+3)^1 = (2a +3)

OpenStudy (anonymous):

but -1 -(-2) =1 not -2 - (-1)

OpenStudy (precal):

I see you are getting this.

OpenStudy (precal):

ok tell me what the book said and I will work it out for you

OpenStudy (anonymous):

ok give me a sec, I'll type it out which will take me a few mins:

OpenStudy (anonymous):

\[The question itss 1/2(a−4)4(2a+3)−2[6(a−4)−(2a+3)] Thats their first step \]

OpenStudy (anonymous):

ops wait something went wrong

OpenStudy (anonymous):

Ok so the question is 3(a−4)5(2a+3)−2−1/2(a−4)4(2a+3)−1

OpenStudy (anonymous):

Their first step is 1/2(a−4)4(2a+3)−2[6(a−4)−(2a+3)]

OpenStudy (anonymous):

then they just expand and isimmpify which I get

OpenStudy (anonymous):

crap those are supposed to be exponents

OpenStudy (anonymous):

the 4 and the -2 are exponents

OpenStudy (anonymous):

klj lol thesis doesn't let me backspace for some reason..

OpenStudy (precal):